# Embed-Search-Align: DNA Sequence Alignment using Transformer models

### v3

@@ 5.2 Task transfer from Chromosome 2
Are these results indicative of DNA-ESA’s adaptability to new genomic sequences, rather than a strict adherence to its training data? This would suggest the model’s learning to solve the sequence alignment task rather than memorizing the genome.
@@
Train Reads 2 (seen in training) 3 (populated) Y (populated)
Chr. 2 Pure 98.0±0.43 98.5±0.39 99.0±0.28 97.5±0.46 98.2±0.41 99.3±0.27
ART 97.9±0.43 97.8±0.43 97.8±0.43 97.9±0.43 97.6±0.44 98.0±0.43
@@
Thermus Aquaticus [All] 100.0 (93.0±0.75 @K=1) 100.0 (95.2±0.63, @K=1)
Acidobacteriota [All] 100.0 (93.2±0.74, @K=1) 100.0 (94.9±0.64, @K=1)
Rattus Norwegicus [Chr. 1] 93.3±0.73 95.4±0.62
Rattus Norwegicus [Chr. 2] 93.2±0.74 96.4±0.55
Pan Troglodytes [Chr. 2A] 94.5±0.67 97.1±0.51
Pan Troglodytes [Chr. 2B] 94.3±0.68 97.2±0.50
@@
Table 3: Cross-Species Task Transfer with DNA-ESA: Training on Human Chr. 2, Testing on Diverse Species: DNA-ESA, trained on human Chr. 2, aligns fragments and reads from different species, including Rattus Norwegicus, Pan Troglodytes, Acidabacteriota, and Thermus Aquaticus. These species, are evaluated for read alignment recall using pure and ART-generated reads. The findings, consistent with Table 1, indicate DNA-ESA’s proficiency in modeling DNA sequence structure, beyond simply memorizing training data. Top-$K=1$ tests are conducted on species achieving 100% recall with $K=50$.
@@
Performance: Details on convergence are in Appendix A. Table 2 shows that the performance on unseen human chromosomes (3, Y) is similar to the performance reported in Table 1, despite only training DNA-ESA on Chr. 2. This suggests DNA-ESA’s ability to generalize sequence alignment across chromosomes with different compositions. High Accuracy, Precision, and F1-scores further confirm the method’s specificity at this high recall. In Table 3, a trend of decreasing recall across species is observed: Human (2, 3, Y) $>$ Chimpanzee (2A, 2B) (Blue) $>$ Rat (1, 2), with chimpanzees likely performing better due to genetic similarities with humans (Ijdo et al., 1991). Even distantly related species like Thermus Aquaticus and Acidobacteriota show significant recall, highlighting DNA-ESA’s task transferability beyond simple data memorization.
5.3 Ablation Studies
We’ve conducted additional experiments on DNA-ESA, detailed in the Appendix: (A) Sec. C.1 covers DNA-ESA’s performance on noisy reads from ART simulators and the PacBio CCS dataset. (B) In Sec. C.2, we explore its performance under various $d_{SW}$ bounds ({5%,10%}) and top-$K$ settings ({10,20}) per chromosome. High performance at lower $K$ values suggests a more effective representation space, while the $d_{SW}$ range assesses fragment retrieval accuracy without exact index matches. (C) Sec. C.3 describes outcomes when omitting diversity priors in searches across the whole genome (top-$K$), resulting in predictably lower performance. (D) Sec. C.4 evaluates DNA-ESA’s effectiveness with different read lengths, including those beyond $Q=250$. Notably, performance improves with longer reads, demonstrating a Zero-shot effect for lengths not included in training.
6 Future Work
@@
While DNA-ESA shows promising recall performance when compared to traditional algorithmic methods, its current computational speed, at 200 reads per minute, is a limiting factor for extensive genomic studies involving millions of reads. Despite this, the alignment task’s inherent parallelizability (outlined in [P1]) offers avenues for efficiency improvements. Our ongoing and future efforts are directed towards overcoming this time efficiency bottleneck, as detailed in Appendix Sec. B. We are considering various optimization strategies, including model compilation to speed up inference and enhanced parallelization in vector store searches and fragment-read alignment. Additionally, we aim to enhance DNA-ESA’s performance with shorter read lengths, exploring alternative training methods, integrating more diverse read/reference features, and employing data augmentation techniques.
@@ 7 Concluding Remarks
Current DNA sequence alignment methods rely on algorithmic approaches that have been refined over decades, incorporating DNA-specific enhancements in both indexing and retrieval. We have introduced an alternate data-driven paradigm. Employing a Transformer-based DNA-ESA encoder, our approach performs sequence alignment through self-supervised learning using contrastive loss. While such methods have previously been used for identifying (approximate) semantic similarity between sequences in written language applications, we demonstrate the surprising ability to find exact overlaps among DNA sequences. Our empirical results also show that the model, once trained, embodies the inherent structure of any DNA sequence (up to context length) irrespective of the location or species origin: a foundation model. DNA-ESA’s innovation lies in its ability to represent various sequence lengths within a shared vector space, facilitating a ‘flat’ search approach. Unlike traditional methods that rely on hierarchical search, DNA-ESA treats shorter reads and longer reference fragments equally, using cosine similarity as a symmetric distance metric. This novel approach not only challenges existing paradigms but also enhances the potential for discovering novel DNA sequence representations. Mirroring the impact of Sentence Transformers in addressing sequence-level tasks in NLP, DNA-ESA, combined with traditional techniques, could lead to breakthroughs like aligning reads to the Pan Genome, where variations in the reference sequences model the genetic variations across individuals.
Appendix
Appendix A Convergence
@@
(a) Trained on the Human Genome
(b) Trained on the Human Chromosome 2
@@
Appendix B Complexity of computing alignment
B.1 Cost of constructing a new representation
@@
The $F^{2}$ complexity follows the basic implementation of attention in transformers, but recent efforts (Beltagy et al., 2020; Kitaev et al., 2020) have developed shortcuts to reduce the cost. These have already been applied to DNA sequence modeling  (Nguyen et al., 2023).
B.2 Vector Store Upstream
@@
B.3 Retrieval Cost
Given a new embedding, ${\epsilon}(G,K)$ is the cost of retrieving top-$K$ nearest neighbors across the fragment embeddings, where $G$ is the length of the reference genome. In modern vector databases, where several hashing techniques such as approximate K-nearest neighbors are used, ${\epsilon}$ scales logarithmically with $G$. This is indeed the key benefit of using such databases instead of a naïve search technique.
B.3.1 Fine-grained Alignment
@@ B.3.2 Total Complexity
$\mathcal{O}(F^{2}+FQK+\epsilon(G,K))\Rightarrow\mathcal{O}(F(F+QK)+\epsilon(G,K))\Rightarrow\mathcal{O}(FQK+\epsilon(G,K))$
Appendix C Ablation studies
C.1 Performance on high-noise reads
In Table 4, we report the performance of DNA-ESA on two noisy datasets: (i) ART-generated (MiSeqv3-based) reads of length $Q=250$ with Phred quality score $Q_{PH}\in[10,20],I,D=1\%$ and (ii) reads from PacBio CCN as generated on the Ashkenazim Trio (Son) (retrieved from the Genome in a Bottle resource (GIAB) Zook et al. (2016)). For (ii), the raw reads are 10 kilobases long and for evaluation we consider random subset of $1000$ reads derived to Chr. 2. Reads specific to a chromosome are filtered using the pbmm2 (Minimap) package Li (2018). Reads input into the aligner are cropped to length $250,350$ and $500$. Note that in these experiments, we take care to increase the fragment-to-fragment overlap $|\mathcal{F}_{i}\bigcap\mathcal{F}_{i+1}|$ from $250$ to $500$. DNA-ESA performance is reported with respect to the Bowtie-2 performance. While Bowtie-2 has a $>99.9\%$ recall, DNA-ESA isn’t far behind and is comparable (within $1\%$) to the results on less noisy reads reported in Table 1. While the recall is lower, the quality of the recovered fragments is comparable: the average Smith-Waterman distance for all successfully aligned {read, fragment} pairs are similar.
MiSeqv3 - Q = 250
(via ART, PH in [10,20], I,D:1%)
PacBio, CCS - pbmm2, chr2
(Ashkenazim Trio, Son)
Model
Very-high
Sensitivity
Alignment
SW (500)
Very-high
Specificity
Alignment
SW (500)
Q=250 Q=350 Q=500
Bowtie-2 >99.9 490.2
87.0 ± 0.94
(–mp 1000
–np 1000)
490.4 <+2.0% <+2.8% <+2.5%
DNA-ESA
94.4 ± 0.60
(dSW=30)
488.3
89.0 ± 0.87
(@dSW=15)
490.6 97.5 ± 0.47 96.5 ± 0.55 97.0 ± 0.51
Table 4: DNA-ESA performance on noisy reads: Noisy reads are simulated using ART ($Q_{PH}=20$, $I,D=1\%$) and derived from an external PacBio CCS read dump available in GIAB. The reads from PacBio are filtered to those derived from Chr. 2 as determined using the pbmm package Li (2018). In this evalaution, DNA-ESA is contextualized with Bowtie-2 Langmead et al. (2018), a popular conventional alignment tool. On both datasets, DNA-ESA performs slightly worse than Bowtie-2. Despite this, the quality of the retrieved fragments – as measured using the SW distance – is very similar. Note that in all cases, the performance of DNA-ESA is comparable (within $1\%$) to those reported in Table. 1.
C.2 Sweeping Top-K and SW distance bound
In Table 5, we report the performance of DNA-ESA with larger Smith-Waterman distance bounds $d_{SW}\in\{10\%,20\%\}$ and a smaller number of recalled fragments $K\in\{10,20\}$. Distance bound $d_{SW}<10\%$ is equivalent to an acceptable mismatch of at most $\sim 8$ bases between the read (length $Q=250$) and recalled reference fragments. We observe that with small $K$, the recall is $>91\%$ (exact index match). As the distance bound increases, performance predictably improves; however, the performance gain from $d_{SW}=25$ to $d_{SW}=50$ is small. This implies that many retrieved fragments are high-quality retrievals; i.e. of high-likelihood to align with the read, and do not benefit from a more generous distance bound to improve recall performance significantly. Furthermore, decreasing the top-$K$ per chrosomosome from $20\rightarrow 10$ does not substantially worsen performance ($\sim 2\%$) indicating that the optimal retrievals are usually the closest in the embedding space.
$Q_{PH}\in[30,60]$ $Q_{PH}\in[60,90]$
$I$ $D$ Exact $d_{SW}<5\%$ $d_{SW}<10\%$ Exact $d_{SW}<5\%$ $d_{SW}<10\%$
Top-10 / chromosome
0.0 0.0 91.5±0.81 96.6±0.54 96.6±0.54 91.9±0.79 96.4±0.55 97.3±0.49
0.01 92.0±0.79 96.5±0.55 96.8±0.52 92.0±0.79 96.2±0.57 97.6±0.47
0.01 0.0 91.9±0.79 96.4±0.55 96.5±0.55 92.0±0.79 96.6±0.54 97.4±0.48
0.01 91.8±0.80 96.3±0.56 97.2±0.50 91.5±0.81 96.3±0.56 97.3±0.49
Top-20 / chromosome
0.0 0.0 93.4±0.73 97.5±0.47 98.0±0.43 93.7±0.71 97.7±0.45 97.8±0.44
0.01 93.2±0.74 97.2±0.49 98.2±0.41 93.6±0.72 97.5±0.47 97.8±0.45
0.01 0.0 93.0±0.74 97.9±0.44 97.9±0.43 92.6±0.76 97.4±0.48 97.8±0.44
0.01 93.0±0.74 97.4±0.48 97.8±0.45 93.4±0.72 97.8±0.45 98.1±0.42
Table 5: Sequence alignment recall of DNA-ESA sweeping top-K and $d_{SW}$: The various parameters are described in Sec. 5.1. DNA-ESA presents a recall of $>97\%$ across several read configurations rivaling contemporary algorithmic models. Performance predictably degrades as more noise is introduced into a read. Performance improves with larger search radius (top-$K$), higher quality reads ($Q_{PH}$) and large distance bound $d_{SW}$.
C.3 Without diversity priors in top-K
@@
$I$ $D$ Exact $d_{SW}<5\%$ $d_{SW}<10\%$ Exact $d_{SW}<5\%$ $d_{SW}<10\%$
Top-100
0.0 0.0 82.6±1.08 92.0±0.79 93.2±0.74 88.5±0.92 91.9±0.79 93.0±0.74
0.01 82.3±1.09 91.8±0.80 92.7±0.76 89.3±0.89 93.2±0.74 93.9±0.70
0.01 0.0 83.3±1.06 92.3±0.77 93.3±0.73 88.7±0.91 92.4±0.77 93.3±0.73
0.01 82.8±1.08 91.6±0.81 92.5±0.76 88.7±0.91 92.4±0.77 93.2±0.73
Top-50
0.0 0.0 81.1±1.11 90.6±0.84 92.0±0.79 86.7±0.97 90.7±0.84 91.9±0.79
0.01 80.5±1.13 90.4±0.85 91.6±0.80 87.4±0.95 91.7±0.80 92.7±0.76
0.01 0.0 81.3±1.11 90.8±0.83 92.3±0.78 86.8±0.97 91.0±0.83 92.2±0.78
0.01 81.2±1.11 90.4±0.85 91.7±0.80 87.1±0.96 91.2±0.82 92.4±0.77
Table 6: Sequence alignment recall of DNA-ESA - without diversity priors: The various parameters are described in Sec. 5.1. DNA-ESA presents a recall of $\approx 90\%$. Similar to the result presented in the main text, performance improves with larger search radius (top-$K$), higher quality reads ($Q_{PH}$) and large distance bound ($d_{SW}$).
C.4 Across read lengths

### v4

@@ 4.2 Task transfer from Chromosome 2
Are these results indicative of DNA-ESA’s adaptability to new genomic sequences rather than strict adherence to its training data? This would suggest the model’s learning to solve the sequence alignment task rather than memorizing the genome.
@@
Train ART Reads 2 (seen in training) 3 (populated) Y (populated)
Chr. 2 $Q_{PH}\in[30,60]$ 99.74 $\pm$ 0.29 99.74 $\pm$ 0.29 $9$$\pm$ 0.01 99.49±0.46 98.87±0.59 99.99±0.01
Chr. 2 $Q_{PH}\in[60,90]$ 99.37 $\pm$ 0.49 99.36 $\pm$ 0.49 $9$$\pm$ 0.01 98.74±0.64 99.12±0.55 99.5±0.43
@@
Thermus Aquaticus [All] 99.9 $\pm$ 0.01 99.9 $\pm$ 0.01
Acidobacteriota [All] 99.9 $\pm$ 0.01 99.9 $\pm$ 0.01
Rattus Norwegicus [Chr. 1] 95.98±1.07 96.48±1.01
Rattus Norwegicus [Chr. 2] 97.99±0.78 96.49±1.01
Pan Troglodytes [Chr. 2A] 98.5±0.69 95.74±1.11
Pan Troglodytes [Chr. 2B] 95.99±1.08 96.25±1.05
@@
Table 4: Cross-Species Task Transfer with DNA-ESA: Training on Human Chr. 2, Testing on Diverse Species: DNA-ESA, trained on human Chr. 2, aligns fragments and reads from different species, including Rattus Norwegicus, Pan Troglodytes, Acidabacteriota, and Thermus Aquaticus. These species, are evaluated for read alignment recall using ART-generated reads. The findings, consistent with Table 1, indicate DNA-ESA’s proficiency in modeling DNA sequence structure, beyond simply memorizing training data.
@@
Performance: Details on convergence are in Appendix 8. Table 3 shows that the performance on unseen human chromosomes (3, Y) is similar to the performance reported in Table 1, despite only training DNA-ESA on Chr. 2. This suggests DNA-ESA’s ability to generalize sequence alignment across chromosomes with different compositions. High Accuracy, Precision, and F1-scores further confirm the method’s specificity at this high recall. In Table 4, a trend of decreasing recall across species is observed: Human (2, 3, Y) $>$ Chimpanzee (2A, 2B) (Blue) $>$ Rat (1, 2), with chimpanzees likely performing better due to genetic similarities with humans (originstory, ). Even distantly related species like Thermus aquaticus and Acidobacteriota show significant recall, highlighting DNA-ESA’s task transferability beyond simple data memorization.
5 Limitations and Future Work
@@
DNA-ESA has the following challenges that we hope to overcome in future generations of the system: the computational speed, at 10K reads per minute, is a limiting factor for extensive genomic studies involving millions of reads. Despite this, the alignment task’s inherent parallelizability offers avenues for efficiency improvements. Our ongoing and future efforts are directed towards overcoming this time efficiency bottleneck, as detailed in Appendix Sec. 9. We are considering various optimization strategies, including model compilation, to speed up inference and enhanced parallelization in vector store searches and fragment-read alignment. Additionally, we aim to enhance DNA-ESA’s performance at shorter read lengths, exploring alternative training methods and integrating more diverse read/reference features.
@@ 6 Concluding Remarks
Current DNA sequence alignment methods, honed over decades, incorporate DNA-specific enhancements in both indexing and retrieval, relying on algorithmic approaches. However, we introduce an alternative data-driven paradigm, employing a Transformer-based DNA-ESA encoder. Our method performs sequence alignment through self-supervised learning, utilizing contrastive loss. While similar methods have previously found application in identifying approximate semantic similarity in written language, our approach surprisingly excels at identifying exact overlaps among DNA sequences. Empirical results underscore the model’s ability to embody the inherent structure of any DNA sequence, irrespective of location or species origin. A distinguishing feature of DNA-ESA is its ability to map both shorter reads and longer reference fragments into the same space by employing cosine similarity as a symmetric distance metric. This complements existing paradigms that have traditionally used a hierarchical and explicit comparison of base distributions.
DNA-ESA representations would find utility in other related tasks. If, for example, one had to align reads to the Pan Genome liao instead of a single reference, one can populate a single vector store with fragments from all $47$ genome assemblies. Then given a read, the nearest fragment to the read can belong to any reference variant. Note that in this case, we avoid having to construct a separate index for each of the references as is usually done in conventional methods.
One can similarly repurpose the representation space – originally intended for Sequence Alignment – to address the complementary task of Genome Assembly. We conclude this paper with a setup that provides initial proofpoints under reasonable simplifying assumptions about read lengths and overlaps among the reads. The resulting framework will be formalized and benchmarked with reads from actual experiments in our future work.
6.1 A setup for de-novo Genome Assembly using DNA-ESA
Consider a set of reads, $R:=\{r_{1},r_{2},\dots,r_{C}\}$ from which we would like to estimate a reference $\mathcal{R}$. The de-novo assembly task is to estimate an ordering of the reads in $R$, $\hat{R}:=(r_{i},\dots,r_{j},\dots,r_{k})$ such that a concatenation (after removing overlaps) is approximately the original reference. In typical setups, all pairs of reads would need to be compared in order to find those that overlap and construct longer chains of the assembly in an incremental fashion. This may be visualized as a complete graph $K_{C}$ with $C$ nodes (reads) and $C^{2}$ edges (constraints). To avoid the quadratic complexity, modern techniques such as Overlap-Layout-Consensus (OLC) or de Bruijn Graph (DBG) approaches assembly have adopted hot-start heuristics such as base overlap and shared K-mer counts to reduce the number of pairwise read comparisons.
The representation space modeled by DNA-ESA admits a comparable assembly setup. Fig. 2 suggested emergent 1D manifolds in the representation space – comprising the embeddings of the reads – $h(r_{1}),h(r_{2}),\dots,h(r_{C})$ – such that the relative positions of fragments along the manifold correlated to the relative locations of the fragments along the reference. Therefore, a pairwise read-read comparison within a smaller radius in the embedding space would be sufficient to find reads that overlap. And more generally, a walk along the 1D manifold would constitute a near-optimal assembly.
Figure 5: Motivating the task of de-novo assembly using the distance adjacency matrix of the UMAP-generated $k$-nearest neighbor ($k=10$) network of reads from Thermus Aquaticus: We assume that reads $r_{1},r_{2},\dots,r_{C}$ belonging to a species are embedded into the DNA-ESA embedding space $h(r_{1}),h(r_{2}),\dots,h(r_{C})$ using a model trained on a known reference genome belonging to a different species. For this illustrative numerical example we ensure that reads are created with the following properties: first, no read is a substring of any other read (efficient algorithms for the case where reads can be substrings of others will be covered in our future work); second every read has a corresponding read that it overlaps with for each end, and finally, the union of the reads cover the ground-truth reference genome. Thus, for every read $r_{i}$ there exist reads $r_{k}$ and $r_{j}$ such that the ordered triplet $(r_{j},r_{i},r_{k})$ constitutes a fragment of the reference. In a perfect embedding scenario, the closest neighbors of $r_{i}$ (i.e. $k=1$ in an $k$-NN search) in the embedding space should be $r_{j}$ ad $r_{k}$. For this ideal case, if a network is constructed where each read (node) is connected to its two nearest neighbors, and the nodes are arranged based on the order in which they appear in the reference chromosome, one would get a bi-diagonal adjacency matrix. When ground truth is not known, the adjacency matrix will be a permuted version of a bi-diagonal matrix, but a complete assembly can still be performed by a greedy algorithm that starts with any node and pieces together the neighboring reads in the embedded space in time that grows linearly in the number of reads. Simulated reads, each of length $500$ and with random overlaps of lengths drawn uniformly $\mathcal{U}([150,250])$ with its neighboring reads at each end, are sampled from $4$ of the chromosomes in Thermus Aquaticus, and are embedded into the DNA-ESA embedding space trained with a reference human genome. The sorted adjacency matrix of the UMAP umap -generated $k$-NN ($k=10$) network is presented above for each chromosome. The dark off-diagonal distance values in the $k$-NN adjacency graph with $k=10$ show that the distance properties of actual embeddings of simulated reads from Thermus Aquaticus are almost coincident with that of the ideal case discussed above. Moreover, just as in the ideal case, by finding an approximate solution to the Traveling Salesman Problem in the unsorted network, one can find a walk (the assembly) that is close to the original reference. The results are presented in Table 5.
We provide an initial computational implementation of this idea: (a) The k-nearest neighbor (kNN) graph $G(R,d)$, representing the distances between close-by reads, is computed using the density-aware corrections provided by the UMAP library umap : graph adjacency matrices for each of the chromosomes is visualized in Fig. 5. For this figure, the several reads are ordered manually as they appear in the assembly in order to highlight the connectivity pattern; the bright off-diagonal values indicate a chain-like structure corresponding to the observed manifold; (b) A Hamiltonian cycle computed through the graph that has the lowest total distance corresponds to an assembly, $\hat{R}$. This involves solving the Traveling Salesman Problem (TSP) – we use the 3/2-approximate Christofides algorithm christofides . After a walk across the reads is generated, a second pass through the walk greedily concatenates adjacent reads $(r_{i},r_{i+1})$ by computing the alignment using the SW distance and removing duplicates and overlaps. In cases where adjacent reads do not have any alignment to one another – the walk isn’t guaranteed to be perfect since the DNA-ESA model is heuristical and the TSP algorithm is approximate – the current assembly part is stowed away as a contig contig , and a new contig is instantiated with the incoming read. The fewer the contigs, the better the manifold in the representation space correlates to the ordering of the true reference.
Assembly is conducted across $4$ of the shorter chromosomes in the Thermus Aquaticus genome. We consider (long) reads of length $Q=500$, which have no indels, and the entire corpus of reads is ensured to have full coverage across each chromosome. Each read has an overlap of $\mathcal{U}([150,250])$ bases with another read in the corpus of reads. QUAST quast metrics are reported in Table 5.
Size (in kb) $\uparrow$ No. of Contigs $\downarrow$ N50 (in kb) $\uparrow$ NGA50 (in kb) $\uparrow$
Thermus Aquaticus NZ_CP010823.1 14.4 1 14.4 14.4
NZ_CP010824.1 16.6 1 16.6 16.6
NZ_CP010825.1 69.9 22 6.0 4.4
NZ_CP010826.1 78.7 15 10.3 10.3
Table 5: DNA-ESA performance on de-novo assembly of the Thermus Aquaticus genome: The solution to the approximate TSP applied on the DNA-ESA read-read network generates a set of contigs that can be evaluated with respect to the original reference using QUAST quast . For the two shorter chrosomomes, the resulting assembly was perfect; the entire reference is encapsulated by one contig. Even in the case with longer chromosomes, the number of contigs are still few and of significant length (as indicated by the high N(G)50 score) suggesting a strong correlation between the distance metric in the embedding space and the ordering of the corresponding reads in the assembly.
N50 refers to the length of a contig (in bases) such that the contigs that exceed this length span 50% of the reference genome. NGA50 is a reference-aware N50 metric. The fewer contigs, large N(GA)50 scores in comparison to the large size of the chromosomes, when considered together, demonstrate that the reads are indeed ordered in a relative position-aware manner in the DNA-ESA embedding space with respect to a reference making the de-novo assembly task viable.
{appendices}
7 Appendix
8 A. Convergence
@@ 7 Appendix
((a)) Trained on the Human Genome
((b)) Trained on the Human Chromosome 2
@@
9 B. Complexity of computing alignment
9.1 Cost of constructing a new representation
@@
The $F^{2}$ complexity follows the basic implementation of attention in transformers, but recent efforts (beltagy_longformer_2020, ; kitaev_reformer_2020, ) have developed shortcuts to reduce the cost. These have already been applied to DNA sequence modeling  (nguyen_hyenadna_2023, ).
9.2 Vector Store Upstream
@@
9.3 Retrieval Cost
Given a new embedding, $\epsilon(G,K)$ is the cost of retrieving top-$K$ nearest neighbors across the fragment embeddings, where $G$ is the length of the reference genome. In modern vector databases, where several hashing techniques such as approximate K-nearest neighbors are used, $\epsilon$ scales logarithmically with $G$. This is indeed the key benefit of using such databases.
9.4 Fine-grained Alignment
@@ 9.5 Total Complexity
$\mathcal{O}(F^{2}+FQK+\epsilon(G,K))\Rightarrow\mathcal{O}(F(F+QK)+\epsilon(G,K))$
$\Rightarrow\mathcal{O}(FQK+\epsilon(G,K))$
10 C. Ablation studies
10.1 Without diversity priors in top-K
@@
$I$ $D$ $d_{SW}<1\%$ $d_{SW}<2\%$ $d_{SW}<5\%$ $d_{SW}<1\%$ $d_{SW}<2\%$ $d_{SW}<5\%$
[…]
0.0 0.0 98.56 $\pm$ 0.67 97.60 $\pm$ 0.85 98.56 $\pm$ 0.67 97.76 $\pm$ 0.82 97.76 $\pm$ 0.82 97.44 $\pm$ 0.88
0.01 0.01 98.8 $\pm$ 0.62 98.4 $\pm$ 0.71 98.8 $\pm$ 0.62 99.2 $\pm$ 0.52 98.4 $\pm$ 0.71 98.4 $\pm$ 0.71
Top-50 global
0.0 0.0 91.6 $\pm$ 1.49 92.4 $\pm$ 1.43 93.8 $\pm$ 1.31 90.4 $\pm$ 1.58 93.8 $\pm$ 1.31 94.8 $\pm$ 1.21
0.01 0.01 89.6 $\pm$ 1.64 91.6 $\pm$ 1.49 96 $\pm$ 1.07 94.4 $\pm$ 1.25 94.4 $\pm$ 1.25 93.2 $\pm$ 1.36
Table 6: Sequence alignment recall of DNA-ESA - without diversity priors: The various parameters are described in Sec. 4. When compared to the main results presented in Table 1, it is evident that the addition of diversity priors results in an improvement of $\sim$ in recall. Similar to the result presented in the main text, performance improves with larger search radius in the vector store (top-$K$), higher quality reads ($Q_{PH}$) and large distance bound ($d_{SW}$).
10.2 Across read lengths